OpenStudy (anonymous):

solve for y. 4y+5/3 = y-10/2 a.y=-4 b.y=-5 c.y=-8

OpenStudy (dan815):

hmmm okay what are u having trouble with

OpenStudy (anonymous):

-8

OpenStudy (nurali):

( 4y + 5 ) / 3 = ( y - 10 ) / 2 8y + 10 = 3y - 30 5y = - 40 y = - 8

OpenStudy (anonymous):

Actually u should have typed the question like given below: (4y+5)/3 = (y-10)/2 8y + 10 = 3y -30 5y = -40 i.e. y = -8